### v1

# Believing without Seeing: Quality Scores for Contextualizing Vision-Language Model Explanations

Abstract
When people query Vision-Language Models (VLMs) but cannot see the accompanying visual context (e.g. for blind and low-vision users), augmenting VLM predictions with natural language explanations can signal which model predictions are reliable. However, prior work has found that explanations can easily convince users that inaccurate VLM predictions are correct. To remedy undesirable overreliance on VLM predictions, we propose evaluating two complementary qualities of VLM-generated explanations via two quality scoring functions. We propose Visual Fidelity, which captures how faithful an explanation is to the visual context, and Contrastiveness, which captures how well the explanation identifies visual details that distinguish the model’s prediction from plausible alternatives. On the A-OKVQA and VizWiz tasks, these quality scoring functions are better calibrated with model correctness than existing explanation qualities. We conduct a user study in which participants have to decide whether a VLM prediction is accurate without viewing its visual context. We observe that showing our quality scores alongside VLM explanations improves participants’ accuracy at predicting VLM correctness by 11.1%, including a 15.4% reduction in the rate of falsely believing incorrect predictions. These findings highlight the utility of explanation quality scores in fostering appropriate reliance on VLM predictions.
1 Introduction
Vision-Language Models (VLMs) are being deployed in applications where users who do not have access to the VLM’s visual context; for example, in assisting blind and low-vision individuals Huh et al. (2024); An et al. (2025); Kim et al. (2025), autonomous multimodal digital agents Koh et al. (2024), and in human-robot collaboration Lukin et al. (2018). However, despite recent advances in VLM capabilities, they often exhibit unreliable behavior, including hallucinating visual details Li et al. (2023); Gunjal et al. (2024) and making overconfident predictions Valdenegro-Toro (2024). In scenarios where users cannot directly observe the visual context of the VLM, it becomes imperative to enable users to accurately trust model outputs. How can we provide adequate context for users to establish appropriate reliance on model predictions and explanations?
Prior work has explored the utility of model explanations to support user decision making Wang & Yin (2021); Bansal et al. (2021). However, natural language explanations can be misleading for inaccurate model predictions Joshi et al. (2023); Chaleshtori et al. (2024); Si et al. (2024a); Sieker et al. (2024). In Figure 1, we consider the question “What period of the day does this photo reflect?” where a VLM (incorrectly) answers “Noon” and generates a plausible explanation referencing cues like “a clock on the building” or “lighting and shadows.” This reasoning may appear highly convincing to a user without access to the image, even though the prediction is incorrect. Users are particularly susceptible to such overreliance on explanations in scenarios with such information asymmetry.
Figure 1: VLM explanations that sound plausible can mislead users. Contextualizing VLM explanations with quality scores may help users know when to rely on VLM outputs, but existing explanation qualities are not calibrated with VLM prediction accuracy. We propose evaluating two new qualities of VLM explanations: Visual Fidelity and Contrastiveness. These qualities are better calibrated with VLM correctness, and also help users make better decisions about when to believe VLM predictions.
Existing explanation metrics Hase et al. (2020); Chen et al. (2023); Liu et al. (2023b) are designed to evaluate explanations generated by text-only LLMs, and we find these are poorly calibrated to VLM prediction accuracy. However, considering the explanation with respect to the accompanying visual context can reveal the reliability of the VLM prediction; for instance, identifying hallucinated details in the explanation (“shadows suggesting that the sun is at its peak”) or that the explanation does not mention the critical detail for identifying the correct answer (the time on the clock).
We propose evaluating two new qualities of VLM-generated explanations: Visual Fidelity—how faithfully a VLM explanation reflects the accompanying visual context—and Contrastiveness—how well the explanation rules out alternative answers. We introduce training-free methods to estimate both these qualities without relying on ground-truth quality annotations.
In experiments with three popular VLMs and two visual reasoning tasks, A-OKVQA Schwenk et al. (2022) and VizWiz Gurari et al. (2018), we find that our Visual Fidelity and Contrastiveness qualities better distinguish correct VLM predictions than do existing evaluation qualities Hase et al. (2020); Chen et al. (2023); Liu et al. (2023b), and are also better calibrated with VLM correctness (§3). We conduct several user studies in which participants evaluated VLM predictions without access to visual inputs (§4). We find that augmenting VLM explanations with our proposed qualities improves participants’ ability to distinguish between correct and incorrect VLM predictions. In particular, showing the product of Visual Fidelity and Contrastiveness as a single quality score leads to an 11.1% absolute improvement in user accuracy, along with a 15.4% absolute reduction in the rate of users falsely believing inaccurate predictions. Finally, we show that presenting explanation qualities via natural language descriptions instead of scores further improves user decision making accuracy (§4.2). Overall, our findings highlight the utility of evaluating the quality of VLM-generated explanations and communicating these qualities to users relying on VLM assistants.
2 Quality Scores for VLM Explanations
Visual reasoning is the task of answering a textual question about an image by drawing inferences from what is visually observed. A vision-language model (VLM) is given an input $x=(I,Q)\in\mathcal{X}$ consisting of an image $I$ and a question $Q$, and produces an answer $a\in\mathcal{A}$, a (closed or open) set of options. Additionally, the VLM generates a natural language explanation $E$ for its prediction $a$. This explanation can be either a chain-of-thought rationale Wei et al. (2023) that the VLM prediction is conditioned on $(IQ\rightarrow Ea)$ or a post hoc justification generated after the VLM prediction $(IQ\rightarrow aE)$ Wiegreffe et al. (2022b).
Algorithm 1 Evaluate Visual Fidelity
0: Image II, question $Q$, model prediction $a$, explanation $E$
0: Visual Fidelity score $S_{\mathrm{VF}}(E)$
[…]
2:for all $q^{V}_{j}\in\mathcal{Q}^{V}$ do
3:  $a^{V}_{j}\leftarrow m_{\texttt{Verif}}(q^{V}_{j},I)$
4:end for
[…]
Algorithm 2 Evaluate Contrastiveness
0: Question $Q$, explanation $E$, set of possible answers $\mathcal{A}$, model prediction $a_{0}\in\mathcal{A}$
0: Contrastiveness score $S_{\mathrm{Contr.}}(E)$
[…]
2:for all $a_{j}\in\mathcal{A}$ do
3:  $h_{j}\leftarrow m_{\texttt{QA}\rightarrow\texttt{S}}(Q,a_{j})$
4:end for
[…]
We define an explanation quality scoring function $S_{\varphi}(E;I,Q,a,\mathcal{A})$ [^1] For brevity, we note arguments for $S$ only when relevant. that maps a VLM-generated explanation $E$ to a score in $[0,1]$ reflecting a specific quality $\varphi$ of the trust cue of an explanation. Existing qualities for natural language explanations, primarily developed and evaluated in text-only settings, typically look at how well an explanation supports the model’s prediction Hase et al. (2020), its informativeness Chen et al. (2023), and its plausibility according to common sense priors Liu et al. (2023b). However, evaluating only these qualities is insufficient for VLM-generated explanations since they do not consider the relationship between the explanation and the visual context. Further, we find that these explanation qualities are poorly calibrated with prediction accuracy (§3) and therefore not useful for a user in determining whether to trust the VLM’s prediction or not.
To address these gaps, we introduce two image-anchored trust cues: Visual Fidelity and Contrastiveness. (Figure 1).
2.1 Visual Fidelity
Visual Fidelity quality captures how faithful the VLM’s explanation is to the contents of the image. The explanation, typically spanning multiple sentences, may mention several details about the image in order to explain its prediction. However, it is possible that some of these details are hallucinated by the model Gunjal et al. (2024). For example, in Figure 1, the VLM justifies its prediction by incorrectly claiming that the shadows indicate that the sun is at its peak. Such misleading details can sway users who cannot observe the VLM’s visual context. Evaluating the Visual Fidelity of an explanation could potentially alert users when important explanation details may be hallucinated.
We measure the visual fidelity of an explanation $E$ with respect to the image $I$ by decomposing the explanation into a set of facts about the image and individually verifying each fact (Algorithm 1). Conceptually, this is an oracle scoring function that would check each atomic claim against ground-truth visual Q&A; we approximate it with an LLM-as-judge, using a verifier $m_{\texttt{Verif}}$ (GPT-4o by default), and we audit this approximation with expert annotations and verifier swaps at Appendix˜F. Concretely, we prompt a question generation LLM $m_{\texttt{QGen}}$ to generate a set of verification questions $\mathcal{Q}^{V}$ that confirm visual details that have been mentioned in the VLM-generated explanation. The model has to generate questions such that answering “yes” confirms the presence of the detail in the image (see Table˜6 for the full LLM prompt). The number of questions generated by $m_{\texttt{QGen}}$ is dependent on the explanation and may vary across instances.
For each verification question $q^{V}_{i}\in\mathcal{Q}^{V}$, a verifier VLM $m_{\texttt{Verif}}$ produces an answer $a^{V}_{i}\in\{\text{yes, no}\}$. The Visual Fidelity score $S_{\mathrm{VF}}(E)$ is calculated as the fraction of questions for which the verifier $m_{\texttt{Verif}}$ answers “yes”:
$S_{\mathrm{VF}}(E)=\frac{\sum_{i=1}^{|\mathcal{Q}^{V}|}\mathbbm{1}\{a^{V}_{i}=\text{``yes''}\}}{|\mathcal{Q}^{V}|}$
A high Visual Fidelity score indicates that the rationale faithfully represents the contents of the image.
2.2 Measuring Contrastiveness
This quality captures whether the VLM’s explanation mentions all relevant visual details to identify the correct answer. The evidences and reasoning in the explanation may not identify the visual details that correctly distinguish the correct answer from alternatives. For instance, the explanation in Figure 1 fails to mention the time on the clock—the key visual cue for correctly answering the question. When such relevant cues are missed, the explanation may inadvertently support other alternative answers apart from the VLM’s prediction. Evaluating the Contrastiveness of an explanation could potentially alert users to when the model explanation fails to identify relevant visual cues.
[…]
We begin by masking mentions of all answers $a\in\mathcal{A}$ in the explanation $E$ to prevent label leakage from affecting the entailment model. Then, for each possible answer $a_{j}\in\mathcal{A}$, a paraphraser LLM $m_{\texttt{QA}\rightarrow\texttt{S}}$ paraphrases the question-answer pair $(Q,a_{j})$ into a declarative sentence $h_{j}$. Finally, we calculate the probability that the masked explanation $P$ entails the hypothesis $h_{j}$ using an entailment model $m_{\texttt{NLI}}$. The Contrastiveness $S_{\mathrm{Contr.}}(E)$ score is computed as the relative entailment probability for the predicted hypothesis $h_{0}$, compared to the entailment probability over all possible hypotheses:
$S_{\mathrm{Contr.}}(E)=\frac{\textsc{Pr}_{\texttt{NLI}}(P\text{ entails }h_{0})}{\sum_{a_{j}\in\mathcal{A}}\textsc{Pr}_{\texttt{NLI}}(P\text{ entails }h_{j})}.$
A high Contrastiveness score indicates that the explanation identifies visual cues that not only support the VLM prediction, but also eliminate plausible alternatives.
2.3 Combining Visual Fidelity and Contrastiveness
We study how complementary our scoring functions are by computing a single quality score that combines the two explanation quality scores $S_{\mathrm{VF}}(E)$ and $S_{\mathrm{Contr.}}(E)$. We evaluate three combinations: the average, product, and minimum of the Visual Fidelity and Contrastiveness quality scores.
$S_{\mathrm{avg}}(E)=\frac{S_{\mathrm{VF}}+S_{\mathrm{Contr.}}}{2};\quad S_{\mathrm{Prod}}(E)=S_{\mathrm{VF}}\times S_{\mathrm{Contr.}};\quad S_{\min}=\min(S_{\mathrm{VF}},S_{\mathrm{Contr.}}).$
3 Are Visual Fidelity and Contrastiveness Indicative of VLM Correctness?
We evaluate our quality scoring functions on their ability to indicate the accuracy of a VLM prediction.
Visual Reasoning Tasks.
We evaluate our explanation qualities on two visual reasoning tasks: A-OKVQA Schwenk et al. (2022)—a multiple-choice VQA benchmark that requires reasoning over images using external knowledge and commonsense—and VizWiz Gurari et al. (2018)—an open-ended VQA task consisting of questions asked by blind and low-vision users about images they captured on their mobile phones. We sample 500 questions from the validation set of each task. Section˜G.2 contains details about the tasks’ pre-processing and evaluation. We do not evaluate Contrastiveness or combined quality scoring functions on VizWiz, which lacks a specified set of possible answers for each question. [^2] We tried turning this into a multiple-choice task by mining negatives from GPT-4o, however we found that the quality of negatives was not very good.
Table 1: Discriminability scores (higher is better) for different Quality Scoring Functions across three models and two visual reasoning tasks. We evaluate significance of Discriminability scores using an unpaired t-test; ∗∗∗ indicates significance at $p<0.001$, ∗∗ indicates $p<0.01$, indicates $p<0.05$.
A-OKVQA VizWiz
LLaVA Qwen2.5 GPT-4o LLaVA Qwen2.5 GPT-4o Average
Simulatability $0.106^{**\phantom{*}}$ $247^{***} 0.0$$98^{\phantom{***}} 0.0$$088^{*\phantom{**}}$ $0.158^{**}$ $130^{*} 0.0$940.094
Informativeness $18^{\phantom{***}} 0.000$$0^{\phantom{***}} 0.0$$69^{\phantom{***}} 0.0$$93^{*\phantom{**}} 0.0$$007^{\phantom{**}}$ $113^{\phantom{*}} 0.0$170.017
Plausibility $043^{**\phantom{*}} 0.0$$22^{\phantom{***}} 0.0$$28^{\phantom{***}} 0.0$$04^{\phantom{***}} 0.0$$20^{\phantom{**}} 0.0$$02^{\phantom{*}} 0.0$310.031
Visual Fidelity $181^{***} 0.0$$80^{***} 0.0$$085^{***}$ $^{***} 0.0420.042^{*\phantom{*}}$ $0.024^{\phantom{*}}$ 0.115
[…]
Models.
We experiment with three popular vision-language models: LLaVA-v1.5-7B Liu et al. (2023a), Qwen2.5-VL-7B Bai et al. (2025), and GPT-4o Hurst et al. (2024) [^3] We use the gpt-4o-2024-05-13 checkpoint.. Answers and rationales are generated in a two-step process: the model is first prompted to predict an answer to the question based on the image, and then prompted to generate a natural language explanation for its prediction. Table˜3 contains the exact prompts we use for answer and explanation generation.
Additionally, our quality scoring functions use several model-based tools. For computing Visual Fidelity, GPT-4o is used both for generating verification questions and answering those questions. For computing Contrastiveness, we use GPT-4o to mask answers in the explanation and paraphrase question-answer pairs into declarative sentences, and the entailment model from Sanyal et al. (2024).
Baseline Explanation Qualities.
We compare our proposed quality measures against three established, text-only explanation qualities, each of which also returns a score in the range $[0,1]$. Simulatability Hase et al. (2020) evaluates whether an explanation offers sufficient evidence to logically justify the model’s prediction. Informativeness Chen et al. (2023) evaluates whether the explanation introduces new information to justify a prediction beyond just re-starting the question and prediction. Commonsense Plausibility Liu et al. (2023b) evaluates whether the explanation is in accordance with commonsense knowledge of everyday situations.
Evaluating Quality Score Calibration.
We evaluate each quality scoring function on its ability to indicate whether a VLM prediction is accurate based on its explanation. Specifically, we consider two evaluation metrics: Discriminability (Disc) and Expected Calibration Error (ECE).
On an evaluation set of $N$ visual reasoning instances, $\text{Disc}(S_{\varphi})$ evaluates a quality scoring function $S_{\varphi}$ by calculating the difference between the mean quality score assigned to instances with accurate predictions ($\text{Acc}(a_{i})=1$) and inaccurate predictions ($\text{Acc}(a_{i})=0$):
$\text{Disc}(S_{q})=\mathbb{E}_{\begin{subarray}{c}1\leq i\leq N\\ \text{Acc}(a_{i})=1\end{subarray}}\bigl[S_{q}(E_{i})\bigr]-\mathbb{E}_{\begin{subarray}{c}1\leq i\leq N\\ \text{Acc}(a_{i})=0\end{subarray}}\bigl[S_{q}(E_{i})\bigr].$
We further calculate whether the difference in means between these two distributions ($S_{\varphi}(E)$, for accurate predictions, versus $S_{\varphi}(E)$, for inaccurate predictions) is significant using an unpaired t-test.
Expected Calibration Error (ECE) Guo et al. (2017) is typically used to evaluate how accurately a model’s confidence estimate reflects the true accuracy of its predictions. We evaluate our quality scoring functions using ECE by interpreting the quality scoring functions as confidence estimates.
Table 2: Expected Calibration Error (ECE, lower is better) for explanation qualities.
A-OKVQA VizWiz
LLaVA Qwen2.5 GPT-4o LLaVA Qwen2.5 GPT-4o Average
Simulatability $0.288$ $0.219$ $0.302$ $0.305$ $0.292$ $0.494$ $0.316$
Informativeness $0.332$ $0.152$ $0.228$ $0.429$ $0.168$ $0.446$ $0.293$
Plausibility $0.162$ $0.270$ $0.317$ $0.110$ $0.310$ $0.325$ $0.248$
Visual Fidelity $0.207$ $0.137$ $0.099$ $0.271$ $0.136$ $0.160$ 0.168
Contrastiveness $0.176$ $0.136$ $0.209$ 0.174
Avg(VF, Contr.) $0.109$ $0.053$ $0.090$ 0.084
Prod(VF, Contr.) $0.164$ $0.154$ $0.233$ 0.183
Min(VF, Contr.) $0.147$ $0.144$ $0.227$ 0.173
Results.
Tables 1 and 2 compare three sets of explanation qualities: existing ones developed for text-only explanations (Simulatability, Informativeness, Plausibility), our proposed qualities (Visual Fidelity, Contrastiveness), and combinations of Visual Fidelity and Contrastiveness.
Figure 2: Calibration curves for various quality scoring functions when evaluating explanations generated by Qwen2.5-VL-7B on the A-OKVQA dataset.
We observe that on A-OKVQA, across all three VLMs, our proposed quality scoring functions achieve higher Discriminability scores and lower Expected Calibration Errors than existing scoring functions. We also see that all Discriminability scores are significant at $p<0.001$, indicating that both our quality scoring functions are capable of distinguishing accurate and inaccurate VLM predictions. Further, combining the Visual Fidelity and Contrastiveness quality scores results in even higher Discriminability scores and lower calibration errors, highlighting complementary utility. On VizWiz, we similarly find that Visual Fidelity is better calibrated than baseline quality scoring functions, and is better at distinguishing correct and incorrect predictions made by LLaVA.
Figure 2 shows calibration curves for the baseline qualities, our proposed qualities, and an average of Visual Fidelity and Contrastiveness, for explanations generated by Qwen2.5-VL-7B on the A-OKVQA task. We observe that Informativeness’s low calibration error is achieved by assigning high scores to all explanations. Our proposed quality scores achieve a lower calibration error while assigning a wider spread of scores across all evaluation samples.
4 Are VLM Explanation Quality Scores Helpful to Users?
Calibration alone does not necessarily translate to downstream utility for users in real-world scenarios Vodrahalli et al. (2022). To what extent can providing $\mathrm{VF}$ and $\mathrm{Contr.}$ scores alongside VLM explanations help users more accurately decide when to believe VLM predictions?
4.1 User Study Setup
We simulate a real-world setting of a user relying on a VLM assistant when the user cannot directly view the VLM’s visual input. Human participants assess the accuracy of VLM predictions.
Figure 3: Our study interface where users are shown the visual reasoning question, VLM prediction and explanation, and optionally one or more quality scores (using simplified language descriptions of both qualities). Here, the user believes that the VLM prediction is incorrect.
For a visual reasoning question $(I,Q)$ from the A-OKVQA and VizWiz tasks, participants are shown the question $Q$, the VLM’s prediction $a$ and explanation $E$, and optionally an explanation quality score $S_{\varphi}(E)$. Importantly, participants cannot see the image $I$ accompanying the question. Participants have to decide whether they believe the model’s answer is correct or incorrect based on the provided context. They can also indicate “unsure” if they feel the information is insufficient for making a reliable judgment. This design helps us measure how presenting different quality scores to a user affects user decision making. Figure 3 contains an example of how information is presented in our user study interface.
Study Details.
We conduct a between-conditions user study, where users makes predictions about VLM correctness based on either none or a subset of the explanation quality scores. We conduct user studies on a subset of 100 questions each from the A-OKVQA and VizWiz tasks, with each subset containing 50 correct and 50 incorrect predictions by LLaVA-v1.5-7B. We select these subsets to approximately match the ECE of $\mathrm{VF}$ and $\mathrm{Contr.}$ scores (if available) with the full evaluation set to maintain a similar distribution of quality scores in our subset.
In a given setting (either none or some of the quality scores are presented to users), each question is annotated by three participants, yielding 300 annotations. Each participant completes 10 questions, also balanced to contain five instances where the model is correct and five where it is incorrect; participants are not informed of this balancing. Participants are randomly assigned to one of the study conditions (control or one of the treatment variants) and do not participate in more than one condition. In total, we collect 300 human annotations per dataset and condition (3 users × 100 questions), capturing user reliance behavior under varying explanation and quality signal conditions.
Evaluation Metrics.
Since participants are asked to judge whether the VLM’s prediction is correct or incorrect, we first exclude responses marked as “unsure.” Following prior work on evaluating human-AI collaboration Joshi et al. (2023); Ma et al. (2024); Srinivasan & Thomason (2025); Wiegreffe et al. (2022a), we then calculate User Accuracy over the remaining responses, and the degree of appropriate reliance using two metrics: Over-Reliance and Under-Reliance. Over-Reliance represents the fraction of interactions where the user believed the VLM prediction was correct, when in fact it was incorrect. Similarly, Under-Reliance is the fraction of interactions where the user mistakenly believed the VLM was incorrect. For both reliance metrics, lower values are preferred.
Figure 4: Effect of showing users different quality scores on User Accuracy, Over-Reliance and Under-Reliance. Error bars represent the standard deviation of the data. Asterisks denote improvements over the explanation‐only baseline using a bootstrap significance test (*: $p<0.05$, **: $p<0.01$).
4.2 RQ1: Does providing explanation quality scores improve user reliance?
We first evaluate whether showing users our proposed quality scores alongside VLM explanations helps users more accurately assess VLM correctness. We compare communicating: 1) only the Visual Fidelity score, 2) only the Contrastiveness score, 3) both scores, and 4) a product of the two scores. Figure 9 contains examples of how these quality scores are communicated to users. We additionally include a control setting where only an explanation is shown without a quality score, and a random setting where users are shown a quality score randomly sampled between 0 and $1$.
Results.
Presenting explanation quality scores, either individually or in combination, consistently improves user performance across all three metrics (Figure 4). Relative to the control explanation-only condition and the Random Score baseline, all treatment conditions lead to higher user accuracy. On A-OKVQA, the most pronounced improvements are observed when showing the product of Visual Fidelity and Contrastiveness, which results in an 11.1% increase in user accuracy, a 15.4% reduction in over-reliance, and a 5.3% reduction in under-reliance. These gains suggest that quality signals help users not only better identify correct model predictions, but also avoid being misled by incorrect ones.
Other treatment variants also show meaningful improvements. Visual Fidelity and Contrastiveness alone both reduce over-reliance by approximately 2–4%, and displaying both scores side-by-side further improves user outcomes. On VizWiz, the trends are similar; addition of quality scores leads to improvements in both accuracy and over-reliance. Together, these findings highlight the utility of explanation quality scores in helping users make more informed and calibrated decisions.
Figure 5: Ablation on presentation types on A-OKVQA dataset: holding the numeric signal fixed (Prod($\mathrm{VF}$, $\mathrm{Contr.}$), $\mathrm{VF}$, or $\mathrm{Contr.}$), we vary the on-screen label (shown as VF, shown as Contr, or simple confidence). User Accuracy ($\uparrow$) and Over-Reliance ($\downarrow$) are effectively unchanged by naming/framing.
4.3 RQ2: Which aspects of the presentation affect user decision accuracy?
Building on RQ1, we explore which presentation choices are significant. We explore two dimensions:
(i) Labeling only (RQ2a): In this scenario, we display the numeric value, while the name displayed on the screen changes (e.g., it could be shown as $\mathrm{VF}$, $\mathrm{Contr.}$, or a generic “AI confidence”).
(ii) Format/content (RQ2b): Here, a scalar score is compared to a descriptive summary that presents evidence (e.g., verified visual claims for $\mathrm{VF}$ or plausible alternatives for $\mathrm{Contr.}$).
4.3.1 RQ2a: Do names and framings of a score affect user reliance?
One concern is whether our observed gains stem from what the score communicates versus how it is labeled to users. In A-OKVQA, using the same underlying score distributions (Prod($\mathrm{VF}$, $\mathrm{Contr.}$), $\mathrm{VF}$, and $\mathrm{Contr.}$), we permuted their on-screen presentation: (i) shown under the correct name, (ii) mislabeled as the other qualities or shown as a generic “AI confidence” without mentioning any quality.
Results.
As shown in Figure 5, renaming/genericizing the numeric signal yields nearly identical outcomes in both User Accuracy and Over-Reliance: differences are small and largely within error bars across all three bases. Thus, we observed that user behavior is driven by the numeric signal’s distribution rather than its label.
Figure 6: Effect of showing numeric and descriptive qualities on User Accuracy and Over-Reliance. Error bars show standard deviation of the data. Asterisks denote improvements between descriptive setting against the corresponding numeric setting with a bootstrap significance test (**: $p<0.01$).
4.4 RQ2b: Do descriptive expressions of quality scores affect user decision accuracy?
Figure 7: Here, the study interface shows the participant descriptive text versions of the Visual Fidelity and Contrastiveness qualities.
Our proposed quality scores are interpretable. For Visual Fidelity, we can observe which visual details in the explanation were and were not verified by the verifier VLM. For Contrastiveness, we can examine which alternative answers were also assigned a high entailment probability. We compare the effect of showing explanation qualities as numeric scores versus interpretable text descriptions on user decision making. Figure 7 shows an example of our proposed qualities as text descriptions.
Results.
As shown in Figure 6, we observe that descriptive formats perform comparably to, and in some cases slightly better than, their numeric counterparts. While showing both quality scores and for $\mathrm{VF}$ individually, the descriptive versions lead to improved user accuracy and reduced over-reliance. However, descriptive treatments hinder performance for $\mathrm{Contr.}$, leading to reduced accuracy. These findings suggest that while numeric scores offer a compact signal, descriptive formats may help reduce over-reliance by making the explanation evaluation process more transparent and interpretable.
4.5 RQ3: How does calibration of quality scores impact user decision making?
Our results in Figure 4 compared quality scores with different levels of calibration, which were also presented with different messages to users. To isolate the relationship between calibration error and user decision accuracy from the presentation of explanation qualities, we run user studies by showing scores from different distributions (Visual Fidelity scores, Contrastiveness scores, and average and product combinations of the two) while presenting the scores to users using the same message (“AI Confidence that the explanation is accurate”).
Results.
In Figure 8, we compare the ECE of different explanation qualities [^4] ECE is computed over the 100 samples used in the user study, not the full evaluation set of 500 samples. and the resulting User Accuracy and Over-Reliance. We observe that there is a negative correlation between calibration error and the downstream user decision making accuracy, with ECE explaining $\approx 60\%$ of the variance in User Accuracy and Over-Reliance respectively. These findings indicate that calibration error of the quality scores is an important determiner of downstream utility to users.
Figure 8: Relationship between ECE of different quality scores and their downstream utility to users.
5 Related Work
Our work builds on insights from evaluation of explanation quality and their utility for decision support, and adapt them for a vision-language setting.
Evaluating Explanation Quality.
Prior work on evaluating explanation quality has largely focused on text-only domains, proposing metrics that assess informativeness Chen et al. (2023), prediction support Wiegreffe et al. (2022b); Hase et al. (2020), acceptability or helpfulness to users Wiegreffe et al. (2022a), and consistency across reasoning chains Golovneva et al. (2023); Prasad et al. (2023). Our work goes beyond textual plausibility by introducing two novel explanation qualities tailored for vision-language settings. Visual Fidelity has been explored for evaluating faithfulness of model-generated image captions Madhyastha et al. (2019); Lee et al. (2023). Our work builds on these ideas to do zero-shot evaluation of explanations.
Explanations for Decision Support.
Most research on utility of natural language explanations for decision support has found that plausible-sounding explanations Jin et al. (2023) often mislead users into overrelying on inaccurate model predictions Joshi et al. (2023); Si et al. (2024b); Chaleshtori et al. (2024); Sieker et al. (2024). Our work adds to these findings by showing that evaluating and communicating explanation qualities to users can reduce overreliance.
6 Conclusions and Future Work
As VLMs are increasingly deployed in settings where users lack access to the model’s visual inputs, users must be able to discern when to rely on model predictions using secondary cues like explanations. However, providing explanations alone can mislead users into believing the model even when it is incorrect. We propose evaluating two unexplored, complementary explanation qualities: Visual Fidelity and Contrastiveness, and also introduce scoring functions for measuring both qualities. We find that our proposed quality scoring functions are well calibrated with model correctness, compared to existing notions of explanation quality. Through several user studies, we also demonstrate real-world utility of these scores, as users presented with these quality scores alongside VLM predictions and explanations are able to improve their task accuracy and reduce over-reliance on the VLM.
Limitations.
A-OKVQA multiple-choice questions may not be representative of real-world visual queries. Further, the design of our Contrastiveness quality limits its application to tasks with a closed-set of possible answers, which is often not available in realistic settings. Finally, we only evaluate on English-language datasets, and conduct user studies with only fluent English speakers.
Future Work.
An important direction is to develop adaptive human-AI reliance strategies that learn when to present explanations, when to suppress them, and when to show only quality scores, depending on the context, task difficulty, or user uncertainty. Additionally, explanation quality scores could be used as training objectives to improve the generation of explanations. Finally, future works should study how explanation quality influences user trust over time in order to better understand how users adapt to model signals and how trust dynamics evolve over repeated human-AI interactions.
Acknowledgments
This research is supported in part by the National Science Foundation (NSF) under grant IIS2403437, the Simons Foundation, and the Allen Institute for AI. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of NSF. B. Joshi was supported by the Apple Scholars in AI/ML PhD Fellowship. This work was partially done when S. Swayamdipta was a visitor at the Simons Institute for the Theory of Computing.
Appendix A Code and Data Availability
Our codebase is available at https://github.com/dill-lab/vlm-quality-scores.
Appendix B Prompts used for Answers and Rationales Generation
We adopt a two‐step, post-hoc justification strategy ($IQ\to aE$) inspired by Wiegreffe et al. (2022b). In the first step, we prompt the model to produce only its answer prediction, without any accompanying rationale. In the second step, we present that predicted answer back to the model and ask it to “Please explain the reasoning behind your answer.”
By separating prediction from explanation, we ensure that the extracted rationale truly reflects the model’s own line of thought under the selected answer. Table˜3 shows the prompts used in every step. Note that the first-step prompt is slightly different for A-OKVQA and VizWiz since the former has answer choices but the latter does not.
Table 3: Prompts used in each step of our two-step method. The first step is to get the model’s predicted answer to the question, and the second step is used to collect the rationale behind the model’s own answer.
Dataset Step System Prompt User Prompt
A-OKVQA Step 1 Answer the question using a single word or phrase from the list of choices. Question: {question}. Choices: {choices}.
Step 2 Please explain the reasoning behind your answer. Question: {question}. Choices: {choices}. The answer is {answer}.
VizWiz Step 1 Answer the user’s question in a single word or phrase. When the provided information is insufficient, respond with ‘Unanswerable’. Whatever the user said, your answer should **always** be a single word or phrase. Question: {question}.
Step 2 Please explain the reasoning behind your answer. Question: {question}. The answer is {answer}.
Appendix C Existing Text-only Qualities Implementations
C.1 Simulatability
To mitigate answer leakage in our Simulatability quality, we first mask all direct mentions of the model’s predicted answer within each rationale by replacing them with the special token <mask>. Given the model’s predicted answer $a$ and its rationale $E$, every occurrence of $a$ in $E$ is substituted with <mask>, yielding the masked rationale $E_{masked}$.
Next, to transform the QA pair into an NLI task, we prompt GPT-4o-mini to convert the original question and its predicted answer into a single declarative hypothesis sentence $H$. Table˜4 shows the model and prompt used to generate $H$. We then feed the masked rationale $E_{masked}$ as the premise and $H$ as the hypothesis into the soumyasanyal/nli-entailment-verifier-xxl model on Hugging Face to obtain an entailment probability $p_{entail}$ as its Simulatability score.
Table 4: Model configuration and prompt used to generate descriptive sentence
Config Assignment
Model gpt-4o-mini-2024-07-18
Max Tokens 1024
Temperature 0.1
User Prompt Integrate the question and the answer into one sentence. For example, given the question "What is the man waiting for?" and the answer "taxi", you should output "The man is waiting for taxi."
Question: {question}
Answer: {answer}
C.2 Informativeness
We utilized GPT-4o to extract the new information contained in the model’s rationale that are not semantically equivalent to the hypothesis.
Table 5: Model configuration and prompt used to evaluate informativeness of a rationale.
Config Assignment
Model gpt-4o-2024-05-13
Max Tokens 1024
Temperature 0.1
User Prompt Please break the following rationale into distinct pieces, and keep only the ones that are not semantically equivalent to the hypothesis. Output the final answer in a Python list format.
Example:
Hypothesis: The man by the bags is waiting for a delivery.
Rationale: The man by the bags is waiting for a delivery, as indicated by the presence of the suitcases and the fact that he is standing on the side of the road. The other options, such as a skateboarder, train, or cab, do not seem to be relevant to the situation depicted in the image.
Output: ["Suitcases are present in the image.", "The man is standing on the side of the road.", "The other options, such as a skateboarder, train, or cab, do not seem to be relevant to the situation depicted in the image."]
Task:
Hypothesis: {hypothesis}
Rationale: {rationale}
Table˜5 shows the model configuration and prompt used to evaluate informativeness of a rationale. After extracting the individual information pieces, we check the size of the resulting list; if it is non–zero, we deem the rationale to be informative.
Appendix D Visual Fidelity
In Visual Fidelity, we use a two-step pipeline to evaluate. Firstly, we generate possible visual verification questions related to the rationale, by providing the question, predicted answer, and rationale to the model, and second, for each question, we provide the question and visual input to the model to ask for verification.
Table 6: Model configuration and prompt used to generate verification visual questions of a rationale.
Config Assignment
Model gpt-4o-2024-08-06
Max Tokens 1024
Temperature 0.1
User prompt: You will be shown a question about an image, along with an answer, and a rationale that explains the answer based on details from the image. Your task is to generate a list of yes/no questions that verify the details about the image that are **explicitly** mentioned in the rationale. Your questions should be phrased such that the answer to that question being yes means that the detail in the rationale is correct. Focus on creating questions that can be visually verified or refuted based on the details provided in the rationale. Ensure the questions are specific and directly pertain to aspects that are visually relevant and mentioned in the rationale. Avoid generating questions about elements that are not mentioned in the rationale, or the rationale explicitly states are not relevant or present. Also avoid generating multiple questions that check for the same visual detail.
Here is one example:
Input:
Question: Why is the person wearing a helmet?
Answer: For safety
Rationale: The person is wearing a helmet because they are riding a bicycle on a busy city street. Helmets are commonly used to protect against head injuries in case of accidents, especially in areas with heavy traffic.
Good Questions:
1. Is the person wearing a helmet while riding a bicycle?
Reason: This question is directly answerable by observing whether the person on the bicycle is wearing a helmet in the image.
2. Is the street in the image busy with traffic?
Reason: This question can be visually verified by looking at the amount of traffic on the street in the image.
Bad Questions:
1. Is the person wearing the helmet because they are concerned about head injuries?
Reason: This question is not good because it assumes the person’s intentions or concerns, which cannot be visually verified from the image.
2. Does wearing a helmet suggest that the person is highly safety-conscious?
Reason: This question relies on inference and external knowledge about the person’s mindset, rather than on observable details from the image.
3. Is there any indication that the person is wearing a helmet for safety reasons?
Reason: This question verifies the answer to the original question, rather than verifying a detail about the image that’s mentioned in the rationale.
4. Is the person wearing a safety vest?
Reason: This question is not good because it tries to verify details about the image that are not explicitly mentioned in the rationale.
5. Is the person not wearing sunglasses?
Reason: This question is not good because it asks for verification by absence and can only be answered with a "no," which is not the preferred type of question.
Respond with a list of (good) questions (without the reasons), starting from ‘1. ’
Table 7: Model configuration and prompt used to verify the visual questions generated from Table˜6.
Config Assignment
Model gpt-4o-2024-08-06
Max Tokens 1024
Temperature 0.1
User prompt: Question: question. Based on the information provided in the image, answer with ‘yes’ or ‘no’. Provide one-word answer only.
Table˜6 refers to the model setting and prompt used for generating the verification questions, while Table˜7 refers to those used for verifying the questions.
Appendix E Example of Model Explanations with Different Qualities
From Table˜8, we show our evaluation framework on four representative VQA instances. For each example, we provide the original question (and answer choices where applicable), image, ground truth answer, the model’s selected answer and rationale, and the corresponding scores for Simulatability, Informativeness, Commonsense Plausibility, Visual Fidelity, and Contrastiveness. Note that for VizWiz, the ground-truth answer is defined as the majority vote among the 10 user annotations from the original dataset, and since VizWiz provides no answer choices, we do not compute a Contrastiveness score.
Table 8: Four VQA examples from both A-OKVQA and VizWiz datasets with model predictions, explanations, and qualities. The predicted answer and explanation are from LLaVA-v1.5-7B model.
Dataset Question Answer Choices Image Ground Truth Predicted Answer LLaVA-v1.5-7B Model Explanation Sim. Info. Plau. VF Contr.
A-OKVQA What is the white substance on top of the cupcakes? butter, mayo, ice cream, icing [Uncaptioned image] icing icing The white substance on top of the cupcakes is icing. This can be inferred from the fact that icing is commonly used as a topping for cupcakes to enhance their flavor and appearance. In the image, there are three cupcakes with icing on top, which indicates that they have been decorated with icing as part of their presentation. Additionally, icing is a common ingredient in cupcake recipes, so it is reasonable to assume that the white substance on top of the cupcakes is indeed icing. 0.391 1.000 0.473 1.000 0.996
A-OKVQA What is the descriptive word for this surface? barren, crowded, minimalist, empty [Uncaptioned image] crowded empty The surface in the image is described as empty. This implies that there are no other objects or items on the surface, making it a minimalist and uncluttered space. The presence of a black cat sitting in front of a computer screen further emphasizes the emptiness of the surface, as the cat is the only object occupying the space. 0.964 1.000 0.566 0.500 0.393
VizWiz What color are those pants? [Uncaptioned image] pink pink The man is wearing pink pajamas, which is an unusual choice for a casual setting. 0.025 1.000 0.397 1.000
VizWiz What is in this bottle? [Uncaptioned image] bacon bits spices The person holding the bottle of spices is likely using it to season their food, adding flavor and aroma to various dishes. Spices are commonly used in cooking as they can enhance the taste of ingredients and create unique flavor combinations. The hand holding the bottle suggests that the person might be preparing a meal or experimenting with different spice blends for their culinary creations. 0.999 1.000 0.566 0.500
Appendix F Reliability and Validity of the Trust Cues
F.1 Validation on Agreement with Human Judgements
To validate the reliability of our automated Visual Fidelity and Contrastiveness qualities, we conducted an internal sanity check on a held‐out sample. We randomly selected 100 examples from our evaluation sets and had a single expert annotator manually assess each instance. The instruction provided to the expert annotator is shown at Table˜11. We binarized our scores at a threshold of $0.5$ (scores $\geq 0.5$ map to $1$, else 0) and compared these labels to the human judgments using Cohen’s $\kappa$:
$\kappa_{\mathrm{VF}}=\,\,0.51\quad\text{and}\quad\kappa_{\mathrm{Contr.}}=\,\,0.44,$
indicating moderate agreement.
F.2 Accuracy of LLM-generated verification questions and answers
To quantify hallucination risk in the VF pipeline, we sampled 50 explanations each from A-OKVQA and VizWiz (total 210 verification questions produced by $m_{\mathrm{QGen}}$ and corresponding answers from $m_{\mathrm{Verif}}$). Three expert annotators judged (i) whether each generated verification question was correctly grounded in the rationale and image, and (ii) whether the verifier’s answer was correct for questions deemed evaluable from the image. We observed that $95.2\%$ of generated questions were correctly grounded and $95.0\%$ of verifier answers were correct. These results support using LLM-as-judge as a cost-efficient approximation to the ideal ground-truth VF scorer.
F.3 Verifier robustness for Visual Fidelity
We use GPT-4o by default for both question generation ($m_{\mathrm{QGen}}$) and verification ($m_{\mathrm{Verif}}$) to isolate pipeline effects. To assess robustness to verifier choice, we swap in smaller open models (Gemma-3n-E4B, Qwen2.5-VL-7B) and report changes in Discriminability (Disc; higher is better) and Expected Calibration Error (ECE; lower is better) on held-out slices.
Table 9: Discriminability of VF scores across verifier models (higher is better). Significance: p<.05p{<}.05, ∗∗ p<.01p{<}.01, ∗∗∗ p<.001p{<}.001.
Dataset Model Accuracy Disc VF(GPT) Disc VF(Gemma) Disc VF(Qwen)
A-OKVQA LLaVA-v1.5-7B $0.696$ $0.181^{***}$ $0.151^{***}$ $0.247^{***}$
A-OKVQA Qwen2.5-VL-7B $0.848$ $080^{***} 0.0$$096^{**\phantom{*}}$ $0.131^{***}$
A-OKVQA GPT-4o $0.910$ $085^{***} 0.0$$044^{\phantom{***}}$ $0.156^{***}$
VizWiz LLaVA-v1.5-7B $0.557$ $240^{***} 0.0$$098^{***}$ $0.193^{***}$
VizWiz Qwen2.5-VL-7B $0.838$ $^{*\phantom{**}} 0.0240.024^{\phantom{***}}$ $0.092^{***}$
VizWiz GPT-4o $0.840$ $024^{\phantom{***}} 0.0$$048^{\phantom{***}}\phantom{-}$ $0.104^{**\phantom{*}}$
Average $0.781$ $109^{\phantom{***}} 0.0$$061^{\phantom{***}}$ $0.154^{\phantom{***}}$
Table 10: Expected Calibration Error (ECE) of VF scores across verifier models (lower is better).
Dataset Model Accuracy ECE VF(GPT) ECE VF(Gemma) ECE VF(Qwen)
A-OKVQA LLaVA-v1.5-7B $0.696$ $0.207$ $0.156$ $0.181$
A-OKVQA Qwen2.5-VL-7B $0.848$ $0.137$ $0.174$ $0.177$
A-OKVQA GPT-4o $0.910$ $0.099$ $0.209$ $0.173$
VizWiz LLaVA-v1.5-7B $0.557$ $0.271$ $0.303$ $0.310$
VizWiz Qwen2.5-VL-7B $0.838$ $0.136$ $0.134$ $0.161$
VizWiz GPT-4o $0.840$ $0.160$ $0.205$ $0.174$
Average $0.781$ $0.168$ $0.197$ $0.196$
As shown in Table˜9 and Table˜10, across A-OKVQA and VizWiz, all three verifiers yield positive discriminability on average (GPT-4o: 0.109; Gemma-3n-E4B: 0.061; Qwen2.5-VL-7B: 0.154), and ECE remains comparably low (0.168/0.197/0.196 respectively). Thus, VF continues to separate correct from incorrect predictions and stays well-calibrated even with using smaller open models as verifiers.
Table 11: Instructions Provided to the Expert Annotator
Criterion Instruction
Visual Fidelity In this task, you’ll view an image, and a question about the image. You will then see an answer to the question given by an AI model, along with an explanation.
Your job is to evaluate whether the details in the explanation are consistent with the image that is shown.
You’ll select one of these choices:
0: The explanation does not mention any details / elements that are directly visible in the image (apart from the prediction itself). Or the explanation mentions details about the image, but one or more of those details are incorrect (they contradict what is visible in the image).
1: The explanation mentions details about the image, and all the details are consistent with the image.
IMPORTANT NOTE: Your job is NOT to check the correctness of the AI model’s answer. It could be that the answer or the logic in the response is incorrect, but the explanation talks about something that is directly in the image. It can also be that the answer is correct, but the explanation does not refer to the image or presents inconsistent details!
IMPORTANT: See examples from this form before proceeding!
Contrastiveness In this task, you’ll view a question about an image (without seeing the image). You will also see an answer from an AI model, along with an explanation (the model has access to the image). Your job is to evaluate whether the explanation meets these two key qualities: You check if the explanation is consistent with the predicted answer. You can ask yourself this question: “Does the explanation provide evidence that matches with the answer it gives?” You check if the explanation covers enough details to reject all other possible answers. You can ask yourself this question: “Does the explanation eliminate all other answers with proper justifications?” By eliminate, we mean that the explanation should provide a strong argument for the selection answer, or strong counter arguments for the other options. You’ll decide whether the explanation has these qualities.
IMPORTANT: See examples from this form before proceeding!
Appendix G User Studies
G.1 Annotator filtering
Table 12: Prolific annotator filtering
Config Criteria
Location United States
Current Country of Residence United States
Primary Language English
Approval Rate 98-100
Number of previous submissions 1000-10000
Highest education level completed Undergraduate degree (BA/BSc/other), Graduate degree (MA/MSc/MPhil/other), Doctorate degree (PhD/other)
Exclude participants from other studies Prohibit any user who took part in a different setting
To ensure high-quality annotations, we recruited annotators on Prolific using the criteria listed in Table˜12. All participants were required to be native English speakers residing in the United States, with approval rates between 98–100% and at least 1000 prior submissions. We further restricted enrollment to individuals holding at least an undergraduate degree. To prevent contamination across experimental conditions, each participant was confined to a single “setting” (i.e., one quality-type configuration).
G.2 Creating the subset splits
For our qualities evaluations from Section˜3, we selected two 500-instance sets. From A-OKVQA we used the first 500 examples from the official validation split. From VizWiz we selected the first 500 validation examples whose majority human annotation was not “unanswerable” and which contained no NSFW content, as filtered by GPT-4o.
To create the subset used for human studies, we create a subset of 100 questions each from the A-OKVQA and VizWiz 500-instance datasets above. These subsets are chosen by sampling 50 times from the validation set; each sampled subset contains randomly sampled 50 correct and 50 incorrect instances from the validation set. On the A-OKVQA dataset, we picked the subset for which the average Expected Calibration Error across Visual Fidelity and Contrastiveness is the lowest in 50 samples; on the VizWiz dataset, we picked the subset for which the Expected Calibration Error for Visual Fidelity is the lowest in our 50 samples, as Contrastiveness is not available on the VizWiz dataset. From Table˜13 this maintains a similar ECE distribution of these quality measures to the full set from Section˜3.
Table 13: Accuracy and Expected Calibration Error (ECE) for different qualities. “Subset selected” rows correspond to the 100-question subsets (50 correct / 50 incorrect) selected and the other two rows represent LLaVA-v1.5-7B on the A-OKVQA and VizWiz 500-instance datasets.
[…]
AOKVQA 0.696 0.207 0.176 0.164 0.147 0.109 0.288 0.332 0.162
AOKVQA (subset selected) 0.500 0.270 0.237 0.133 0.130 0.227 0.372 0.490 0.075
VizWiz 0.557 0.271 0.305 0.429 0.110
VizWiz (subset selected) 0.500 0.236 0.379 0.390 0.071
G.3 Attention Incentives
To ensure that participants engaged carefully with each instance’s annotation, we combined a per-item timer with small monetary bonuses and penalties.
G.3.1 Timer Implementation
To ensure that participants spent sufficient time on each stage (and did not simply skim and click through), we imposed a per-item timer in all of our human studies. The user may only start their selection after the timer ends. For each question, we computed the explanation’s “reading time” as
$\mathrm{reading\_time}=\frac{\#\text{words in explanation}}{238\ \text{words/minute}}$
where 238 wpm is the average adult reading speed Brysbaert (2019).
We then capped the total display time at
$\mathrm{reading\_time}+10\ \text{seconds}$
to cover the question, answers, and any qualites shown.
G.3.2 Bonus Payments
To further motivate careful reading and discourage guessing, we tied each response to a small bonus bank: correct answers earned $\mathdollar 0.10$; incorrect answers incurred $\mathdollar 0.10$ (with the bonus bank floored at $\mathdollar 0$ so it would not harm their base payments); selecting “I’m unsure based on the information provided” resulted in no change. Participants were paid their accumulated bonus in addition to the base participation fee of $\mathdollar 2$ for annotating 10 instances.
G.4 User Studies Settings & Examples
Table 14: Summary of the 14 one‐step human‐study settings.
Setting Description
Show Explanation Only No quality scores shown
Random Score Random scoring baseline (uniform random distribution from $[0,1]$), shown as a simple confidence score “AI Confidence that the explanation is correct”.
[…]
AVG(VF, Contr) Average of VF and Contrast scores, shown as a simple confidence score “AI Confidence that the explanation is correct”.
VF num Show Visual Fidelity numeric score “AI Confidence that the explanation accurately describes the image details”
VF desc Show at most two descriptive sentences converted from visual questions which are verified by verifier VLM $m_{\texttt{Verif}}$, and at most two from questions which are not verified by $m_{\texttt{Verif}}$
Contr. num Show Contrastiveness score “AI Confidence that the explanation rules out the other choices”
Contr. desc Show the other answer options $a_{j}\neq a_{0}$ s.t. $\textsc{Pr}_{\texttt{NLI}}(P\text{ entails }h_{j})\geq 0.5$
Both Numeric Display both VF num and Contr. num messages
Both Descriptive Display both VF desc and Contr. desc messages
VF shown as Conf VF score displayed as a simple confidence score “AI Confidence that the explanation is correct”
Contr shown as Conf Contr. score displayed as a simple confidence score “AI Confidence that the explanation is correct”
[…]
(a) Control (no quality)
(b) VF – Numeric quality
(c) Contr. – Numeric quality
[…]
(e) Both VF and Contr. – Numeric quality
(f) VF – Descriptive quality
(g) Contr. – Descriptive quality
(h) Both VF and Contr. – Descriptive quality
Figure 9: Explanation quality messages for each instruction condition. Subfigure 9(a) is the baseline with no quality displayed; the next three (9(b)9(e)) show numeric qualities; the remaining four (9(f)9(h)) show descriptive ones.
Table 15: User-study results on A-OKVQA and VizWiz in the one stage setting. Each row received 300 annotations (100 questions $\times$ 3 annotations per question).
Dataset Setting Unsure Rate Not Unsure Accept Rate Not Unsure Accuracy False Accept Rate False Reject Rate
A-OKVQA Control 8.7% ± 1.6% 76.3% ± 2.6% 59.1% ± 3.0% 60.7% ± 2.7% 14.0% ± 1.5%
Random Score 14.3% ± 2.0% 70.0% ± 2.9% 59.9% ± 3.1% 50.7% ± 2.5% 18.0% ± 1.7%
[…]
AVG(VF, Contr.) 7.3% ± 1.5% 74.8% ± 2.6% 66.2% ± 2.8% 52.7% ± 2.5% 10.0% ± 1.3%
VF num 7.3% ± 1.5% 76.3% ± 2.6% 62.9% ± 2.9% 58.7% ± 2.6% 10.0% ± 1.3%
VF desc 9.7% ± 1.7% 69.4% ± 2.8% 68.6% ± 2.8% 45.3% ± 2.4% 11.3% ± 1.3%
Contr. num 9.0% ± 1.7% 77.7% ± 2.5% 64.1% ± 2.9% 56.7% ± 2.6% 8.7% ± 1.2%
Contr. desc 6.0% ± 1.4% 71.6% ± 2.7% 61.3% ± 2.9% 56.7% ± 2.6% 16.0% ± 1.6%
Both Numeric 10.7% ± 1.8% 72.8% ± 2.7% 66.4% ± 2.9% 48.7% ± 2.5% 11.3% ± 1.3%
Both Descriptive 8.7% ± 1.6% 68.2% ± 2.8% 67.2% ± 2.8% 46.0% ± 2.4% 14.0% ± 1.5%
VF shown as Conf 7.0% ± 1.5% 77.8% ± 2.5% 61.3% ± 2.9% 60.0% ± 2.7% 12.0% ± 1.4%
Contr. shown as Conf 8.3% ± 1.6% 74.5% ± 2.6% 69.1% ± 2.8% 50.0% ± 2.5% 6.7% ± 1.0%
Prod shown as VF 7.7% ± 1.5% 65.7% ± 2.9% 67.5% ± 2.8% 44.4% ± 2.4% 16.0% ± 1.6%
Prod shown as Contr. 8.0% ± 1.6% 70.3% ± 2.8% 68.5% ± 2.8% 44.7% ± 2.4% 13.3% ± 1.4%
VizWiz Control 7.3% ± 1.5% 70.5% ± 2.7% 59.4% ± 3.0% 55.3% ± 2.6% 20.0% ± 1.7%
VF num 9.3% ± 1.7% 65.1% ± 2.9% 64.7% ± 2.9% 44.7% ± 2.4% 19.3% ± 1.7%
VF desc 10.3% ± 1.8% 67.3% ± 2.9% 69.1% ± 2.8% 42.0% ± 2.4% 13.3% ± 1.4%
In the main user study, we present 14 settings in total. Table˜14 summarizes all 14 different settings and Figure˜9 shows examples of all 8 different categories of messages presented to users. Table˜15 contains the main user-study results of user behavior patterns.
G.5 Supplementary Human Studies
In addition to the study described in §4.2 and Section˜G.4, we ran a supplementary user study to evaluate how explanations and qualities affect users’ trust in model predictions. Each of the 10 questions was presented in three successive stages; after each stage, participants indicated whether they believed the model’s answer was correct or were unsure:
  1. 1.
    Answer Only: Participants viewed only the question, answer choices (if available), and model prediction.
  2. 2.
    With Explanation: Participants were provided with AI-generated rationales alongside predictions.
  3. 3.
    With Explanation + Quality: Qualities (varied from our experiment settings, e.g. Visual Fidelity and Contrastiveness) were displayed alongside explanations.
This three-stage design of the user study enables us to track how users’ confidence in the model’s correctness evolves as they receive additional information.
Timed Stages in Supplementary Human Studies
To standardize attention across conditions, we also enforced a stage-specific timer, where the users can only make their selections at a stage after the timer at that stage ends:
  1. 1.
    Answer Only: fixed 5 seconds
  2. 2.
    With Explanation: explanation reading time (words / 238 wpm) (roughly 10–40 seconds)
  3. 3.
    Explanation + Quality: fixed 5 seconds
Bonus Payments in Supplementary Human Studies
Participants were paid a \$2 base fee and could earn up to \$1 in performance‐based bonuses, which were awarded only during Stage 3 (Explanation + Quality; see Section˜G.3.2).
Table 16: User‐study results on A‑OKVQA and VizWiz in the three stages setting.
Dataset VLM User study setting #Ann. After Stage 1 After Stage 2 After Stage 3
Unsure Rate User Accuracy Unsure Rate User Accuracy Unsure Rate User Accuracy
A‑OKVQA LLaVA-v1.5-7B VF (numeric) 300 $65.3\%\pm 2.8\%$ $67.3\%\pm 4.6\%$ $19.0\%\pm 2.3\%$ $63.4\%\pm 3.1\%$ $5.0\%\pm 1.3\%$ $63.5\%\pm 2.9\%$
Contr (numeric) 300 $65.7\%\pm 2.7\%$ $58.3\%\pm 4.9\%$ $17.7\%\pm 2.2\%$ $60.7\%\pm 3.1\%$ $6.3\%\pm 1.4\%$ $61.2\%\pm 2.9\%$
Both VF and Contr (numeric) 300 $77.3\%\pm 2.4\%$ $58.8\%\pm 6.0\%$ $38.0\%\pm 2.8\%$ $54.8\%\pm 3.7\%$ $9.7\%\pm 1.7\%$ $64.2\%\pm 2.9\%$
Avg(VF, Contr) 300 $86.0\%\pm 2.0\%$ $54.8\%\pm 7.8\%$ $35.3\%\pm 2.8\%$ $63.9\%\pm 3.5\%$ $6.0\%\pm 1.4\%$ $63.5\%\pm 2.9\%$
VF (descriptive) 300 $67.3\%\pm 2.7\%$ $59.2\%\pm 5.0\%$ $19.7\%\pm 2.3\%$ $59.3\%\pm 3.2\%$ $3.7\%\pm 1.1\%$ $62.3\%\pm 2.9\%$
Contr (descriptive) 300 $63.0\%\pm 2.8\%$ $64.9\%\pm 4.6\%$ $17.0\%\pm 2.2\%$ $63.1\%\pm 3.1\%$ $8.7\%\pm 1.6\%$ $59.5\%\pm 3.0\%$
Both VF and Contr (descriptive) 290 $70.7\%\pm 2.7\%$ $47.1\%\pm 5.4\%$ $23.8\%\pm 2.5\%$ $57.9\%\pm 3.3\%$ $8.3\%\pm 1.6\%$ $64.7\%\pm 2.9\%$
VizWiz Qwen2.5-VL-7B VF (numeric) 300 $57.7\%\pm 2.9\%$ $59.8\%\pm 4.4\%$ $5.7\%\pm 1.3\%$ $65.7\%\pm 2.8\%$ $1.7\%\pm 0.7\%$ $65.1\%\pm 2.8\%$
VF (descriptive) 300 $58.7\%\pm 2.8\%$ $66.1\%\pm 4.3\%$ $17.7\%\pm 2.2\%$ $65.2\%\pm 3.0\%$ $4.7\%\pm 1.2\%$ $67.5\%\pm 2.8\%$
Table˜16 shows that as users progress from seeing only the model’s answer to viewing explanations and then explanations with quality scores, their unsure rate steadily decreases while user accuracy correspondingly increases, with the largest gains observed when descriptive qualities are provided alongside explanations. These findings support our findings that richer, more interpretable quality signals can meaningfully improve users’ trust calibration.
Appendix H Computational Resources Spent & Total Cost
The complete breakdown of all monetary expenditures across the study is given in Table˜17. In brief, the computational expenses include six hours of A100 GPU time for running the vision‐language models, plus OpenAI API calls for generating predicted answers, explanations, and computing qualities. Human evaluation costs cover two Prolific studies—a one‐stage study as shown in Section˜G.4 and a three‐stage follow‐up from Section˜G.5 (including up to \$1.00 in per‐participation bonuses; see Section˜G.3.2).
Altogether, these sum to an overall expense of approximately \$2,827.
Table 17: Monetary Cost Breakdown
Study & Step Activity Total Cost
Predicted answer & Explanation Generation (LLaVA-v1.5-7B & Qwen2.5-VL-7B) GPU inference (A100, 6h)
Predicted answer & Explanation Generation (GPT-4o) OpenAI API call $\approx$\$30
Qualities Computation OpenAI API call $\approx$\$120
Human Study (One-stage) Prolific participant pay \$1,561
Supplementary Human Study (Three-stages) Prolific participant pay \$1,116
$\approx$\$2,827